# Ensemble-Based Unsupervised Discontinuous Constituency Parsing by Tree Averaging

### v1

@@ 4 Related Work
Constituency parsing carries a long history in natural language processing research (Charniak, 2000; Klein, 2005; Kallmeyer and Maier, 2010; Li et al., 2019). Different setups of the task have been introduced and explored, including supervised and unsupervised, continuous and discontinuous constituency parsing (Shen et al., 2018a, b, 2019; Fernández-González and Gómez-Rodríguez, 2020, 2021, 2023; Yang et al., 2023). In the unsupervised setup, researchers typically define a joint distribution over the parse structure and an observable variable, e.g., the sentence itself, and maximize the observable variable’s likelihood through marginalization (Kim et al., 2019a, b). To the best of our knowledge, previous unsupervised parsing studies are all restricted to binary structures to squeeze the marginalization space, and we are the first to address non-binary unsupervised parsing and show non-binarity is beneficial to parsing performance.
Ensemble methods strategically combine multiple models to improve performance, rooted in the bagging concept where different data portions are used to train multiple models  (Breiman, 1996a; Hastie et al., 2009). To build an ensemble, straightforward methods include averaging and voting (Breiman, 1996a, b). For outputs with internal structures, minimum Bayes risk decoding (MBR; Bickel and Doksum, 2015) can be used to build an ensemble, where the vote is the negative risk in MBR. However, existing MBR approaches are mostly selective, where the output is selected from a candidate set (Kumar and Byrne, 2004; Titov and Henderson, 2006; Shi et al., 2022). Shayegh et al. (2023) develop a generative MBR that searches for a binary continuous tree. This work extends the Shayegh et al. (2023) and searches in the space of discontinuous constituency trees, leading to significant algorithmic design and theoretical analysis.
@@ 5 Conclusion
In this work, we address ensemble-based unsupervised discontinuous constituency parsing by tree averaging, where we provide comprehensive complexity analysis and develop an efficient search algorithm to obtain the average tree. Our experiments on Dutch and German demonstrate the effectiveness of our ensemble unsupervised method. To the best of our knowledge, we are also the first to address, and show the importance of, non-binary structures in unsupervised constituency parsing.
@@ 6 Limitations
First, our work is rooted in a parsing project. We provide a series of theoretical analyses and algorithmic designs for averaging constituency trees under different setups. Following the trajectory, our ensemble approach may be extended to trees or graphs beyond the parsing domain. We are happy to explore this direction as future work.
@@
Third, our theoretical analysis leaves an open problem about averaging binary trees. However, pointing to open problems is usually considered a contribution (instead of a weakness) in theoretical computer science. Our theoretical analysis is also crucial to the understanding of our algorithms, because we now know that our proposed method works at the level of Open Problem 1.
@@ A.1 Proof of Theorem 1
Following the spirit of Shayegh et al. (2023), we may design a DP table with two axes being the start and end of every component of a constituent, thus $2F$-axes for fan-out $F$. For outputs being non-binary, our DP algorithm requires one additional axis to indicate the number of nodes in a subtree. [^10] If the output is restricted to binary trees, we do not need to track the number of nodes in a subtree, because it is always $12n-$ 1 for $n$ leaves.
@@
For recursion, we divide a constituent $c$ into smaller sub-constituents based on breaking points $\bm{j}=(0,j_{1},j_{2},\cdots,j_{2F-1},|c|)$, where $|c|$ is the number of words in $c$, shown in Figure 6. We further join sub-constituents based on the parity (even or odd) of their indexes.
@@
where $c[b:e]$ denotes a constituent containing $b$th to $e$th words in $c$, and $\oplus$ symbol is the joint of constituents. We also distribute the capacity of the number of nodes into $e_{\bm{j}}$ and $o_{\bm{j}}$ branches by setting $s$ as the $o_{\bm{j}}$’s share and assigning the rest to $e_{\bm{j}}$.
To find the best breaking points and shares, we iterate over all possible values for them:
@@
according to Eqn. (2). Problem 1, as a decision problem, can then be solved trivially by checking whether the score in (14) is greater than or equal to $z$.
@@ A.2 Proof of Theorem 2
In §A.1, we provide a DP algorithm and by backtracking we can easily obtain, for any $\tau$, the constituency tree $T$ that maximizes $\sum_{c\in T}\mathpzc h(c)$ such that $|C(T)|=\tau$ (i.e., the tree having $\tau$ nodes). The DP algorithm, in fact, works for any scoring function $\mathpzc h$ over constituents. Therefore, we have a polynomial-time solver for the following problem:
Fixed-Size Maximization. Given a scoring function $\mathpzc h$ over constituents of a length-$n$ sentence and a natural number $\tau$ such that $n<\tau<2n$, what is the tree $T$ that maximizes $\sum_{c\in T}\mathpzc h(c)$ with $|C(T)|=\tau$?
@@
$=\operatorname*{argmax}_{T:|T|=\tau}\sum_{c\in C(T)}\underbrace{\sum_{k=1}^{K}\frac{\mathds{1}[c\in C(T_{k}(s))]}{\tau+|C(T_{k})|}}_{\hat{\mathpzc h}(c)}$ (17)
@@ A.3 Proof of Theorem 3
For reduction, we use our problem to solve the max clique problem with $G=\langle V,E\rangle$ and $z$ being given.
Without loss of generality, we may assume that $G$ does not contain any vertex that is connected to all other vertices for the rest of this proof. This is because such a vertex is guaranteed to appear in any maximal clique and can be pruned. Then the original problem is equivalent to determining whether the pruned graph has $(z-a)$-many vertices, where $a$ is the number of pruned vertices. Notice that such pruning can be accomplished in polynomial time, thus not changing the complexity category.
@@
We next show that Problem 3 given $\{T_{i}\}_{i:u_{i}\in{V}}$ and the value $\tilde{z}=\frac{2{z}+2|{V}|(n+1)}{{z}+n+2}$ is equivalent to the max clique problem given the graph $G$ and the number ${z}$.
@@
$=\frac{2|C(T)|+2|{V}|(n+1)}{|C({T})|+n+2}$ (22)
where $|C({T})|$ is the same as the number of vertices in the corresponding clique. This shows a clique tree ${T}$ with a score of $\tilde{z}=\frac{2{z}+2|{V}|(n+1)}{{z}+n+2}$ corresponds to a clique with ${z}$ vertices.
@@
[Case 1] If the answer to Problem 3 is “no,” it means that there does not exist a tree (either a clique tree or a general tree) $T$ satisfying $\sum_{u_{i}\in{V}}F_{1}({T},{T}_{i})\geq\tilde{z}$. Thus, there does not exist a clique $Q$ satisfying $|Q|\geq z$.
[Case 2] If the answer to Problem 3 is “yes,” there must exist a tree $T$ satisfying the condition $\sum_{u_{i}\in{V}}F_{1}({T},{T}_{i})\geq\tilde{z}$. Notice that such a tree is not guaranteed to be a clique tree, but in this case, we can show there must also exist a clique tree satisfying the condition.
@@
$\tilde{z}$ $\leq\sum_{k=1}^{K}F_{1}(T,T_{k})$ (23)
$=\sum_{k=1}^{K}\frac{2\sum_{c\in C(T)}\mathpzc h(c)}{|C(T)|+|C(T_{k})|}$ (24)
$\leq\sum_{k=1}^{K}\frac{2\sum_{c\in C(T)\backslash C_{0}(T)}\mathpzc h(c)}{|C(T)|-|C_{0}(T)|+|C(T_{k})|}$ (25)
$=\sum_{k=1}^{K}\frac{2\sum_{c\in C(T^{\prime})}\mathpzc h(c)}{|C(T^{\prime})|+|C(T_{k})|}$ (26)
$=\sum_{k=1}^{K}F_{1}(T^{\prime},T_{k})$ (27)
@@
Since $T^{\prime}$ is a clique tree that satisfies $\sum_{k=1}^{K}F_{1}(T^{\prime},T_{k})\geq\tilde{z}$, we find the solution to the max clique problem is also “yes.”
@@ A.4 Proof of Lemma 1
A selection of constituents corresponds to a constituency tree over a sentence if and only if (1) it includes all the single words and the whole-sentence constituent, referred to as trivial constituents, and (2) all of its constituents are compatible with each other (already satisfied by clique the definition).
@@ A.5 Proof of Theorem 5
Consider any constituent $c_{0}\in C(T^{*})\backslash P$. We construct $T^{\prime}$ by removing $c_{0}$ from $T^{*}$, i.e., $C(T^{*})=C(T^{\prime})\cup\{c_{0}\}$. Based on Eqn. (2), we have
@@ A.6 Proof of Theorem 6
[Part (b)] We assume by way of contradiction that a constituent $c_{0}$ with a hit count of $K$ does not appear in the average tree $T^{*}$. As shown in (a), $c_{0}$ is compatible with every constituent in $T^{*}$; hence, there exists a constituency tree $\hat{T}$ such that $C(\hat{T})=C(T^{*})\cup\{c_{0}\}$. Based on Eqn. (2), we have
@@ Appendix D Case Studies
Consider the last three words “moeten worden opgesteld” (must be established) [^11] We provide English interpretations by ChatGPT-4 with the prompt: “Translate the following Dutch sentence. Give word-by-word literal translation.’’. We see the ensemble output follows the majority votes (three out of five, namely Individuals 22224444, and $5$) and detects the correct structure of this phrase. Moreover, the ensemble is able to detect the discontinuous constituent “in sommige gevallen … niet toereikend zijn” (in some cases … not sufficient be), suggested by Individuals $4$ and $5$. On the other hand, we also notice that every individual produces incorrect discontinuous constituents, but our ensemble is able to smooth out such noise and achieve high $F_{1}$ scores.

### v2

@@ 4 Related Work
Constituency parsing carries a long history in natural language processing research (Charniak, 2000; Klein, 2005; Kallmeyer and Maier, 2010; Li et al., 2019). Different setups of the task have been introduced and explored, including supervised and unsupervised, continuous and discontinuous constituency parsing (Shen et al., 2018a, b, 2019; Corro, 2020; Fernández-González and Gómez-Rodríguez, 2020, 2021, 2023; Chen and Komachi, 2023; Yang et al., 2023; Yang and Tu, 2023). In the unsupervised setup, researchers typically define a joint distribution over the parse structure and an observable variable, e.g., the sentence itself, and maximize the observable variable’s likelihood through marginalization (Kim et al., 2019a, b). To the best of our knowledge, previous unsupervised parsing studies are all restricted to binary structures to squeeze the marginalization space, and we are the first to address non-binary unsupervised parsing and show non-binarity is beneficial to parsing performance.
Ensemble methods strategically combine multiple models to improve performance, rooted in the bagging concept where different data portions are used to train multiple models  (Breiman, 1996a; Hastie et al., 2009). To build an ensemble, straightforward methods include averaging and voting (Breiman, 1996a, b). For outputs with internal structures, minimum Bayes risk decoding (MBR; Bickel and Doksum, 2015) can be used to build an ensemble, where the vote is the negative risk in MBR. However, existing MBR approaches are mostly selective, where the output is selected from a candidate set (Kumar and Byrne, 2004; Titov and Henderson, 2006; Shi et al., 2022). Petrov and Klein (2007) formulate MBR for supervised constituency parsing and propose to search for the global optimum decoding when the risk allows for dynamic programming. Smith and Smith (2007) extend the idea to non-projective dependency parsing. In our previous work Shayegh et al. (2024), we formally introduce generative MBR and develop an algorithm that searches for a binary continuous tree. This paper extends our previous work and searches in the space of discontinuous constituency trees, leading to significant algorithmic design and theoretical analysis.
@@ 5 Conclusion
In this work, we address ensemble-based unsupervised discontinuous constituency parsing by tree averaging, where we provide comprehensive complexity analysis and develop an efficient search algorithm to obtain the average tree. Our experiments on Dutch and German demonstrate the effectiveness of our ensemble method. To the best of our knowledge, we are also the first to address, and show the importance of, non-binary structures in unsupervised constituency parsing.
@@ 6 Limitations
First, our work is focused on syntactic parsing, and we have provided a series of theoretical analyses and algorithmic designs for averaging constituency trees under different setups. Following the trajectory, our ensemble approach may be extended to other structures beyond parsing. We are happy to explore this direction as future work. For example, a concurrent study of ours develops algorithms for ensemble-based text generation Wen et al. (2024).
@@
Third, our theoretical analysis leaves an open problem about averaging binary trees. However, pointing to open problems is usually considered a contribution (instead of a weakness) in theoretical computer science. Our theoretical analysis is also crucial to the understanding of our algorithms, because we now know that our proposed method works at the level of Open Problem 1 and can be easily extended to the level of Problem 3 following the technique proposed in Appendix A.2.
Acknowledgments
We would like to thank all reviewers and chairs for their valuable and constructive comments. The research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Amii Fellow Program, the Canada CIFAR AI Chair Program, an Alberta Innovates Program, a donation from DeepMind, and the Digital Research Alliance of Canada (alliancecan.ca). We also thank Yara Kamkar for providing advice on the algorithms.
@@ A.1 Proof of Theorem 1
Following our previous work (Shayegh et al., 2024), we may design a DP table with two axes being the start and end of every component of a constituent, thus $2F$-axes for fan-out $F$. For outputs being non-binary, our DP algorithm requires one additional axis to indicate the number of nodes in a subtree. [^12] If the output is restricted to binary trees, we do not need to track the number of nodes in a subtree, because it is always $12n-$ 1 for $n$ leaves.
@@
For recursion, we divide a constituent $c$ into smaller sub-constituents based on alternation points $\bm{j}=(0,j_{1},j_{2},\cdots,j_{2F-1},|c|)$, where $|c|$ is the number of words in $c$, shown in Figure 6. We further join sub-constituents based on the parity (even or odd) of their indexes.
@@
where $c[b:e]$ denotes a sub-constituent containing $b$th to $e$th words in $c$, and $\oplus$ symbol is the joint of sub-constituents. We also distribute the capacity of the number of nodes into $e_{\bm{j}}$ and $o_{\bm{j}}$ branches by setting $s$ as the $o_{\bm{j}}$’s share and assigning the rest to $e_{\bm{j}}$.
To find the best alternation points and shares, we iterate over all possible values for them:
@@
according to Eqn. (2.2). Problem 1, as a decision problem, can then be solved trivially by checking whether the score in (7) is greater than or equal to $z$.
@@ A.2 Proof of Theorem 2
In §A.1, we provide a DP algorithm and by backtracking we can easily obtain, for any $\tau$, the constituency tree $T$ that maximizes $\sum_{c\in C(T)}\mathpzc h(c)$ such that $|C(T)|=\tau$ (i.e., the tree having $\tau$ nodes). The DP algorithm, in fact, works for any scoring function $\mathpzc h$ over constituents. Therefore, we have a polynomial-time solver for the following problem:
{adjustwidth}
Fixed-Size Maximization. Given a scoring function $\mathpzc h$ over constituents of a length-$n$ sentence and a natural number $\tau$ such that $n<\tau<2n$, what is the tree $T$ that maximizes $\sum_{c\in C(T)}\mathpzc h(c)$ with $|C(T)|=\tau$?
@@
$=\operatorname*{argmax}_{T:|T|=\tau}\sum_{c\in C(T)}\underbrace{\sum_{k=1}^{K}\frac{\mathds{1}[c\in C(T_{k})]}{\tau+|C(T_{k})|}}_{\hat{\mathpzc h}(c)}$ (10)
@@ A.3 Proof of Theorem 3
For reduction, we use a solver for Problem 3 to solve the max clique problem with $G=\langle V,E\rangle$ and $z$ being given.
Without loss of generality, we may assume that $G$ does not contain any vertex that is connected to all other vertices for the rest of this proof. This is because such a vertex is guaranteed to appear in any maximal clique and can be pruned. Then the original problem is equivalent to determining whether the pruned graph has a clique with $(z-a)$-many vertices, where $a$ is the number of pruned vertices. Notice that such pruning can be accomplished in polynomial time, thus not changing the complexity category.
@@
We next show that Problem 3 given $\{T_{i}\}_{i:u_{i}\in{V}}$ and the value $\tilde{z}=\frac{2{z}+2|{V}|(n+1)}{{z}+2n+3}$ is equivalent to the max clique problem given the graph $G$ and the number ${z}$.
@@
$=\frac{2(|C(T)|-n-1)+2|{V}|(n+1)}{|C({T})|+n+2}$ (15)
where $|C({T})|-n-1$ is the same as the number of vertices in the corresponding clique. This shows a clique tree ${T}$ with a score of $\tilde{z}=\frac{2{z}+2|{V}|(n+1)}{{z}+2n+3}$ corresponds to a clique with ${z}$ vertices.
@@
[Case 1] If the answer to Problem 3 is “no,” it means that there does not exist a tree (either a clique tree or a general tree) $T$ satisfying $\sum_{i:u_{i}\in{V}}F_{1}({T},{T}_{i})\geq\tilde{z}$. Thus, there does not exist a clique $Q$ satisfying $|Q|\geq z$.
[Case 2] If the answer to Problem 3 is “yes,” there must exist a tree $T$ satisfying the condition $\sum_{i:u_{i}\in{V}}F_{1}({T},{T}_{i})\geq\tilde{z}$. Notice that such a tree is not guaranteed to be a clique tree, but in this case, we can show there must also exist a clique tree satisfying the condition.
@@
$\tilde{z}$ $\leq\sum_{i:u_{i}\in{V}}F_{1}(T,T_{i})$ (16)
$=\sum_{i:u_{i}\in{V}}\frac{2\sum_{c\in C(T)}\mathds{1}[c\in C(T_{i})]}{|C(T)|+|C(T_{i})|}$ (17)
$\leq\sum_{i:u_{i}\in{V}}\frac{2\sum_{c\in C(T)\backslash C_{0}(T)}\mathds{1}[c\in C(T_{i})]}{|C(T)|-|C_{0}(T)|+|C(T_{i})|}$ (18)
$=\sum_{i:u_{i}\in{V}}\frac{2\sum_{c\in C(T^{\prime})}\mathds{1}[c\in C(T_{i})]}{|C(T^{\prime})|+|C(T_{i})|}$ (19)
$=\sum_{i:u_{i}\in{V}}F_{1}(T^{\prime},T_{i})$ (20)
@@
Since $T^{\prime}$ is a clique tree that satisfies $\sum_{i:u_{i}\in{V}}F_{1}(T^{\prime},T_{i})\geq\tilde{z}$, we find the solution to the max clique problem is also “yes.”
@@ A.4 Proof of Lemma 1
A selection of constituents corresponds to a constituency tree over a sentence if and only if (1) it includes all the single words and the whole-sentence constituent, referred to as trivial constituents, and (2) all of its constituents are compatible with each other (already satisfied by the clique definition).
@@ A.5 Proof of Theorem 5
Consider any constituent $c_{0}\in C(T^{*})\backslash P$. We construct $T^{\prime}$ by removing $c_{0}$ from $T^{*}$, i.e., $C(T^{*})=C(T^{\prime})\cup\{c_{0}\}$. Based on Eqn. (2.2), we have
@@ A.6 Proof of Theorem 6
[Part (b)] We assume by way of contradiction that a constituent $c_{0}$ with a hit count of $K$ does not appear in the average tree $T^{*}$. As shown in (a), $c_{0}$ is compatible with every constituent in $T^{*}$; hence, there exists a constituency tree $\hat{T}$ such that $C(\hat{T})=C(T^{*})\cup\{c_{0}\}$. Based on Eqn. (2.2), we have
@@ Appendix D Case Studies
Consider the last three words “moeten worden opgesteld” (must be established). [^13] We provide English interpretations by ChatGPT-4 with the prompt: “Translate the following Dutch sentence. Give word-by-word literal translation.’’ We see the ensemble output follows the majority votes (three out of five, namely Individuals 22224444, and $5$) and detects the correct structure of this phrase. Moreover, the ensemble is able to detect the discontinuous constituent “in sommige gevallen … niet toereikend zijn” (in some cases … not sufficient be), suggested by Individuals $4$ and $5$. On the other hand, we also notice that every individual produces incorrect discontinuous constituents, but our ensemble is able to smooth out such noise and achieve high $F_{1}$ scores.